# On the conditioning of histopolationL. Bruni Bruno, S. Serra-Capizzano

### v1

@@ 3.1 An inversion formula for the Chebyshev segments
The columns of $V_{d}^{-1}$ are the coefficients that turn a given basis into the Lagrange one, i.e. the only (up to permutations) basis $\{\ell_{1},\ldots,\ell_{d}\}$ for $\mathscr{P}_{d-1}(I)$ such that $\int_{s_{i}}\ell_{j}(x)\mathrm{d}x=\delta_{i,j}$, see [6, Proposition 1]. An explicit formula for the inverse of $V_{d}$ is crucial, as such bases in histopolation are known only for segments in the classes (C1) and (C3).
@@
$\left[V_{d}^{-1}\right]_{i,j}=\left[W_{d}^{-1}\right]_{i,i}\left[V_{d}\right]_{i,j}=\frac{i^{2}}{2d\sin^{2}(i\rho)}\cdot\frac{2\sin(i\tau_{j})\sin(i\rho)}{i}=\frac{i\sin(i\tau_{j})}{d\sin(i\rho)}.$
@@ 3.2 A formula for the conditioning of Chebyshev segments
An immediate consequence of Proposition 3.1 is that the singular values of $V_{d}$ are the diagonal entries of $V_{d}^{\top}V_{d}$. A direct inspection on (10) also shows that the diagonal elements of $V_{d}^{\top}V_{d}$ are decreasingly ordered as, for $i>j$,
$\left[V_{d}^{\top}V_{d}\right]_{i,i}=\frac{2d}{i^{2}}\sin^{2}(i\rho)\geq\frac{2d}{j^{2}}\sin^{2}(j\rho)=\left[V_{d}^{\top}V_{d}\right]_{j,j}.$
This gives the simplified formula
(15) $\kappa_{2}\left(V_{d}\right)=\frac{\sigma_{1}(V_{d})}{\sigma_{d}(V_{d})}=\sqrt{\frac{\lambda_{1}(V_{d}^{\top}V_{d})}{\lambda_{d}(V_{d}^{\top}V_{d})}}=\sqrt{\frac{[V_{d}^{\top}V_{d}]_{1,1}}{[V_{d}^{\top}V_{d}]_{d,d}}}=\sqrt{\frac{d^{2}\sin^{2}(\rho)}{\sin^{2}(d\rho)}}=\left|\frac{d\sin(\rho)}{\sin(d\rho)}\right|$
for the conditioning of $V_{d}$, that hence depends on $\rho$. It is however quite natural to inversely relate $\rho$ and $d$, as done in [4], for instance in such a way that the arcs overlap in their endpoints, to obtain segments that belong to both classes (C1) and (C2). More generally, one has the following result.
@@
Let $\rho=\frac{a}{d}$. Then,
$\lim_{d\to\infty}\kappa_{2}\left(V_{d}\right)=\left|\frac{a}{\sin(a)}\right|.$
@@
In particular, for Chebyshev segments belonging to the intersection of classes (C1) and (C2), the conditioning $\kappa_{2}(V_{d})$ is monotonically increasing towards
$\lim_{d\to\infty}\kappa_{2}\left(V_{d}\right)=\pi/2.$
Proof 3.9.
Plugging the condition $\rho=\frac{a}{d}$ into (15), one gets that
$\kappa_{2}\left(V_{d}\right)=\left|\frac{d\sin(\rho)}{\sin(d\rho)}\right|=\left|\frac{d\sin(\frac{a}{d})}{\sin(d\frac{a}{d})}\right|=\frac{1}{|\sin(a)|}\left|d\sin\left(\frac{a}{d}\right)\right|.$
By considering the limit $d\to\infty$, and setting $t=a/d$, one has
$\lim_{d\to\infty}\kappa_{2}\left(V_{d}\right)=\lim_{d\to\infty}\frac{1}{|\sin(a)|}\left|d\sin\left(\frac{a}{d}\right)\right|=\lim_{t\to 0}\frac{|a|}{|\sin(a)|}\left|\frac{\sin\left(t\right)}{t}\right|=\frac{|a|}{|\sin(a)|}.$
The specification for $a=\pi/2$ is immediately obtained by substituting the value. To show the monotonicity of the conditioning, we shall prove that
$\frac{(d+1)\sin\left(\frac{\pi}{2(d+1)}\right)}{\sin\left((d+1)\frac{\pi}{2(d+1)}\right)}=\frac{(d+1)\sin\left(\frac{\pi}{2(d+1)}\right)}{\sin\left(\frac{\pi}{2}\right)}\geq\frac{d\sin\left(\frac{\pi}{2d}\right)}{\sin\left(\frac{\pi}{2}\right)}=\frac{d\sin\left(\frac{\pi}{2d}\right)}{\sin\left(d\frac{\pi}{2d}\right)}$
@@
for each $d\geq 1$. This can be done by showing that the function $f(x):=x\sin(\pi/(2x))$ is increasing for $x>1$, i.e., that $f^{\prime}(x)>0$ for $x>1$. After differentiation, making the change of variable $t=\pi/(2x)$, it is readily seen that this is equivalent to proving that $g(t):=\tan(t)-t>0$ for $t\in(0,\pi/2)$. The first derivative of this function is $g^{\prime}(t)=\tan^{2}(t)$, which is clearly defined and positive in such an interval, and since $\lim_{t\to 0^{+}}g(t)=0^{+}$, this concludes the proof.
A numerical verification of the case $a=\pi/2$ of Theorem 3.8 is provided in Figure 7.
@@
Figure 7: Bounded conditioning of the Vandermonde matrix in the Chebyshev basis for the class (C2). The parameter here is $a=\pi/2$, so the resulting segments also belong to the class (C1).

### v2

@@ 3.1 An inversion formula for the Chebyshev segments
The columns of $V_{d}^{-1}$ are the coefficients that turn a given basis into the Lagrange one, i.e. the only (up to permutations) basis $\{\ell_{1},\ldots,\ell_{d}\}$ for $\mathscr{P}_{d-1}(I)$ such that $\int_{s_{i}}\ell_{j}(x)\mathrm{d}x=\delta_{i,j}$, see [8, Proposition 1]. An explicit formula for the inverse of $V_{d}$ is crucial, as such bases in histopolation are known only for segments in the classes (C1) and (C3).
@@
$\left[V_{d}^{-1}\right]_{i,j}=\left[W_{d}^{-1}\right]_{i,i}\left[V_{d}\right]_{i,j}=\frac{i^{2}}{2d\sin^{2}(i\rho)}\cdot\frac{2\sin(i\tau_{j})\sin(i\rho)}{i}=\frac{i\sin(i\tau_{j})}{d\sin(i\rho)}$
for $i\neq d$, and
$\left[V_{d}^{-1}\right]_{d,j}=\left[W_{d}^{-1}\right]_{d,d}\left[V_{d}\right]_{d,j}=\frac{d^{2}}{4d\sin^{2}(d\rho)}\cdot\frac{2\sin(d\tau_{j})\sin(d\rho)}{d}=\frac{\sin(d\tau_{j})}{2\sin(d\rho)}$
for $i=d$ (i.e., the same formula as above, with the corrective factor $2$).
@@ 3.2 A formula for the conditioning of Chebyshev segments
An immediate consequence of Proposition 3.1 is that the singular values of $V_{d}$ are the diagonal entries of $V_{d}^{\top}V_{d}$. A direct inspection on (9) shows that the ordering of the diagonal elements of $V_{d}^{\top}V_{d}$ depends on the relationship between $\rho$ and $d$. Nevertheless, monotonicity conditions can be explicitly stated.
Lemma 3.6.
Let $0<\rho<\frac{\pi}{d}$. Then for $1\leq i<j\leq d-1$, one has
$\left[V_{d}^{\top}V_{d}\right]_{i,i}=\frac{2d}{i^{2}}\sin^{2}(i\rho)>\frac{2d}{j^{2}}\sin^{2}(j\rho)=\left[V_{d}^{\top}V_{d}\right]_{j,j}.$
Proof.
Consider the function $f(x)=\sin^{2}(x\rho)/x^{2}$, for $x>0$. One computes that
$f^{\prime}(x)=\frac{2\sin(\rho x)\left(\rho x\cos(\rho x)-\sin(\rho x)\right)}{x^{3}},$
whence $f^{\prime}(x)<0$ for $0<x<\pi/\rho$. This implies that, for $i,j=1,\ldots,d-1$ and $i<j$,
$\frac{\sin^{2}(i\rho)}{i^{2}}>\frac{\sin^{2}(j\rho)}{j^{2}},$
which immediately yields the claim.∎
Remark 3.7.
Although the connection between $\rho$ and $d$ may seem a technical condition, such an inverse relationship between $\rho$ and $d$ is in fact quite natural, as discussed in [6]. Indeed, taking $\rho=\frac{\pi}{2d}$, arcs overlap in their endpoints, and one obtains segments that belong to both classes (C1) and (C2). Any smaller $\rho>0$ gives non-overlapping segments, while $\rho=\frac{\pi}{d}$ results in a loss of unisolvence, as evident from (8). From now on we hence assume that the arc-length parameter is of the form $\rho=a\frac{\pi}{d}$, with $0<a<1$. This choice is consistent with the literature [7] and ensures the well-posedness of the histopolation problem [6].
Lemma 3.6 states that the all diagonal terms of $V_{d}^{\top}V_{d}$, except the last one, are decreasingly ordered, provided that $\rho$ is sufficiently small. Since $V_{d}^{\top}V_{d}$ is a diagonal matrix, the ratio between its largest and smallest element gives the Euclidean conditioning of $V_{d}$. Hence, to obtain an explicit formula for $\kappa_{2}(V_{d})$, we only need to understand how the last diagonal term of $W_{d}:=V_{d}^{\top}V_{d}$ relates with the others.
Lemma 3.8.
Let $\frac{\pi}{2d}\leq\rho<\frac{\pi}{d}$ (i.e., $1/2\leq a<1$). One has
$W_{d}]_{1,1}=2d\sin^{2}(\rho)>\frac{4}{d}\sin^{2}(d\rho)=[W_{d}]_{d,d}$ (19)
Let $0<\rho\leq\frac{\pi}{2d}$ (i.e., $0<a\leq 1/2$). One has
$W_{d}]_{d,d}=\frac{4}{d}\sin^{2}(d\rho)>\frac{2d}{(d-1)^{2}}\sin^{2}((d-1)\rho)=[W_{d}]_{d-1,d-1}.$ (20)
Further, for any $0<\rho<\frac{\pi}{d}$ (i.e., for any $0<a<1$), one has
$\lim_{d\to\infty}\frac{[W_{d}]_{d,d}}{[W_{d}]_{d-1,d-1}}=2.$ (21)
Proof.
The claims for (19) and (20) can be obtained by an analogous analysis to that of Lemma 3.6. To obtain (21), we plug the explicit characterizations given in (10) and compute
$\lim_{d\to\infty}\frac{[W_{d}]_{d,d}}{[W_{d}]_{d-1,d-1}}$ $=\lim_{d\to\infty}\frac{\frac{4d}{d^{2}}\sin^{2}(\rho d)}{\frac{2d}{(d-1)^{2}}\sin^{2}((d-1)\rho)}=\lim_{d\to\infty}\frac{2(d-1)^{2}\sin^{2}(a\frac{\pi}{d}d)}{d^{2}\sin^{2}((d-1)a\frac{\pi}{d})}$
$=2\lim_{d\to\infty}\frac{(d-1)^{2}}{d^{2}}\frac{\sin^{2}(a\pi)}{\sin^{2}(a\frac{d-1}{d}\pi)}=2.$
Observe that the above computation is meaningful, as $0<a\frac{d-1}{d}\pi<a\pi<\pi$, so the denominator never vanishes. ∎
Lemma 3.8 has two important consequences. First, it evidences the limiting role of the case $\rho=\pi/(2d)$ even beyond the context of interpolation discussed in Remark 3.7. By a continuity argument, for arc-lengths around such a value, merging (19) and (20) one deduces the simple formula
$\kappa_{2}\left(V_{d}\right)$ $=\frac{\sigma_{1}(V_{d})}{\sigma_{d}(V_{d})}=\sqrt{\frac{\lambda_{1}(V_{d}^{\top}V_{d})}{\lambda_{d}(V_{d}^{\top}V_{d})}}=\sqrt{\frac{[V_{d}^{\top}V_{d}]_{1,1}}{[V_{d}^{\top}V_{d}]_{d-1,d-1}}}=\sqrt{\frac{(d-1)^{2}\sin^{2}(\rho)}{\sin^{2}((d-1)\rho)}}$ (22)
$=\frac{(d-1)\sin(\rho)}{\sin((d-1)\rho)}$
for the conditioning of $V_{d}$. Note the abosence of the absolute value in the last term, which has been removed due to the positivity of the sine function for $0<\rho=a\pi/d<a(d-1)/d\pi=(d-1)\rho<\pi$ with $0<a<1$. Further, as $\kappa_{2}(V_{d})$ can be obviously lower bounded by the ratio of any two diagonal terms of $W_{d}$, for large values of $d$ we may take the (d1)(d-1)-th and the $d$-th to obtain
$\kappa_{2}(V_{d})\geq\sqrt{\frac{[V_{d}^{\top}V_{d}]_{d,d}}{[V_{d}^{\top}V_{d}]_{d-1,d-1}}},$
and owing to (21) one also gets the asymptotic lower bound
$\lim_{d\to\infty}\kappa_{2}(V_{d})\geq\sqrt{2}.$ (23)
@@
For the Chebyshev segments belonging to the intersection of classes (C1) and (C2), i.e. for $\rho=a\frac{\pi}{d}$ with $a=1/2$, the conditioning $\kappa_{2}(V_{d})$ is monotonically increasing towards
$\lim_{d\to\infty}\kappa_{2}\left(V_{d}\right)=\frac{\pi}{2}.$
Proof.
Plugging the condition $\rho=\frac{\pi}{2d}$ into (22), one gets that
$\kappa_{2}\left(V_{d}\right)=\frac{(d-1)\sin(\rho)}{\sin((d-1)\rho)}=\frac{(d-1)\sin(\frac{\pi}{2d})}{\sin((d-1)\frac{\pi}{2d})}.$
By considering the limit $d\to\infty$, and setting $t=\pi/(2d)$, one has
$\lim_{d\to\infty}\kappa_{2}\left(V_{d}\right)=\lim_{d\to\infty}\frac{(d-1)\sin(\frac{\pi}{2d})}{\sin((d-1)\frac{\pi}{2d})}=\frac{1}{\sin(\frac{\pi}{2})}\lim_{t\to 0}\frac{\left(\pi-2t\right)\sin(t)}{2t}=\frac{\pi}{2}.$
To show the monotonicity of the conditioning, we shall prove that
$\frac{d\sin\left(\frac{\pi}{2d}\right)}{\sin\left(d\frac{\pi}{2d}\right)}=\frac{d\sin\left(\frac{\pi}{2d}\right)}{\sin\left(\frac{\pi}{2}\right)}\geq\frac{(d-1)\sin\left(\frac{\pi}{2(d-1)}\right)}{\sin\left(\frac{\pi}{2}\right)}=\frac{(d-1)\sin\left(\frac{\pi}{2(d-1)}\right)}{\sin\left((d-1)\frac{\pi}{2(d-1)}\right)}$
@@
for each $d\geq 1$. This can be done by showing that the function $f(x):=x\sin(\pi/(2x))$ is increasing for $x>1$, i.e., that $f^{\prime}(x)>0$ for $x>1$. After differentiation, making the change of variable $t=\pi/(2x)$, it is readily seen that this is equivalent to proving that $g(t):=\tan(t)-t>0$ for $t\in(0,\pi/2)$. The first derivative of this function is $g^{\prime}(t)=\tan^{2}(t)$, which is clearly defined and positive in such an interval, and since $\lim_{t\to 0^{+}}g(t)=0^{+}$, this concludes the proof. ∎
A numerical verification of the case $a=1/2$ of Theorem 3.9 is provided in Figure 7.
@@
Figure 7: Bounded conditioning of the Vandermonde matrix in the Chebyshev basis for the class (C2). The parameter is $a=1/2$, so $\rho=\pi/(2d)$. The resulting segments also belong to the class (C1).
3.2.1 Asymptotics for $a$ far from $1/2$
Lemma 3.8, together with an immediate continuity argument, showed that the conditioning of $V_{d}$ around $a=1/2$ is described by the square root of the ratio between $W_{d}]_{1,1}$ and $W_{d}]_{d-1,d-1}$. For small perturbations, hence, Theorem 3.9 provides a good approximation of the limit value. However, if one considers $a$ either close to 0 or $1$, it is possible to reverse the inequalities (19) and (20).
Example 3.10.
Consider $a=1/10$ and $d=3$. One has
$W_{3}]_{3,3}=\frac{4}{3}\sin^{2}\left(\frac{\pi}{10}\right)\approx 0.1274>0.0656\approx 6\sin^{2}\left(\frac{\pi}{30}\right)=[W_{3}]_{1,1}.$
Consider $a=9/10$ and $d=3$. One has
$W_{3}]_{3,3}=\frac{4}{3}\sin^{2}\left(\frac{9}{10}\pi\right)\approx 0.1274<1.8090\approx 2\sin^{2}\left(\frac{3}{5}\pi\right)=[W_{3}]_{2,2}.$
From the perspective of an asymptotic analysis, Example 3.10 does not constitute a real obstacle, as we are interested in the relationship between the elements of $W_{d}$ for large values of $d$. To this end, we gather the results of Section 3.2 in the following Proposition, which is an immediate consequence of Lemma 3.6 and Lemma 3.8.
Proposition 3.11.
Let $0<a<1$. One has
$\kappa_{2}^{2}(V_{d})=\begin{cases}\frac{[W_{d}]_{1,1}}{[W_{d}]_{d,d}}&\text{if }[W_{d}]_{d,d}<[W_{d}]_{d-1,d-1},\\ \frac{[W_{d}]_{d,d}}{[W_{d}]_{d-1,d-1}}&\text{if }[W_{d}]_{d,d}\geq[W_{d}]_{1,1},\\ \frac{[W_{d}]_{1,1}}{[W_{d}]_{d-1,d-1}}&\text{if }[W_{d}]_{1,1}\geq[W_{d}]_{d,d}\geq[W_{d}]_{d-1,d-1}.\end{cases}$ (24)
It remains to establish which case one falls in. First of all, recall (21) from Lemma 3.8. Such a formula implies that, for large values of $d$, the first case of (24) is empty. We now show that, for large values of $d$, there exists a thresholding value $a^{*}$ for $0<a<1$ which separates the remaining two situations in Proposition 3.11.
Proposition 3.12.
There exists $a^{*}$, with $0<a^{*}<1$, such that for each $0<a<a^{*}$ one has
$\lim_{d\to\infty}\frac{[W_{d}]_{d,d}}{[W_{d}]_{1,1}}>1.$ (25)
Proof.
To establish (25) one computes, applying the substitution $t=a\pi/d$,
$\lim_{d\to\infty}\frac{[W_{d}]_{d,d}}{[W_{d}]_{1,1}}$ $=\lim_{d\to\infty}\frac{\frac{4}{d}\sin^{2}(da\frac{\pi}{d})}{2d\sin^{2}(a\frac{\pi}{d})}=\lim_{d\to\infty}\frac{\sin^{2}(a\pi)}{\frac{d^{2}}{2}\sin^{2}(a\frac{\pi}{d})}=2\sin^{2}(a\pi)\lim_{d\to\infty}\frac{1}{d^{2}\sin^{2}\left(a\frac{\pi}{d}\right)}$
$=2\sin^{2}(a\pi)\lim_{t\to 0}\frac{t^{2}(t)}{\sin^{2}(t)}\frac{1}{a^{2}\pi^{2}}=\frac{2\sin^{2}(a\pi)}{a^{2}\pi^{2}}.$
By computing an approximate solution of $\frac{2\sin^{2}(a\pi)}{a^{2}\pi^{2}}>1$, one obtains $a^{*}\approx 0.40$. ∎
For large $d$, one may hence make the claim of Proposition 3.11 more precise.
Theorem 3.13.
One has
$\lim_{d\to\infty}\kappa_{2}(V_{d})=\begin{cases}\frac{a\pi}{\sin(a\pi)}&\text{if }a^{*}<a<1,\\ \sqrt{2}&\text{if }0<a<a^{*}.\end{cases}$ (26)
Proof.
Suppose first that $a\in(0,a^{*})$. Then, by (25), for large $d$ one has $W_{d}]_{d,d}>[W_{d}]_{1,1}$. Due to Lemma 3.6, one in general has $W_{d}]_{1,1}>[W_{d}]_{i,i}$ for any $i=2,\ldots,d-1$, hence
$\lim_{d\to\infty}\kappa_{2}(V_{d})=\lim_{d\to\infty}\sqrt{\frac{[W_{d}]_{d,d}}{[W_{d}]_{d-1,d-1}}}=\sqrt{2},$
as computed in (21).
Suppose then that $a\in(a^{*},1)$. By (25), for large $d$ one has that in such an interval $W_{d}]_{1,1}>[W_{d}]_{d,d}$. Further, by (21), we deduce that in this limit $W_{d}]_{d,d}>[W_{d}]_{d-1,d-1}$. It follows that
$\lim_{d\to\infty}\kappa_{2}(V_{d})=\lim_{d\to\infty}\sqrt{\frac{[W_{d}]_{1,1}}{[W_{d}]_{d-1,d-1}}}=\lim_{d\to\infty}\sqrt{\frac{2d\sin^{2}\left(a\frac{\pi}{d}\right)}{\frac{2d}{(d-1)^{2}}\sin^{2}\left(a\pi\frac{(d-1)}{d}\right)}}$
and, by applying the same substitution $t=a\pi/d$ and reasoning of Proposition 3.11, one readily computes that
$\lim_{d\to\infty}\kappa_{2}(V_{d})=\frac{a\pi}{\sin(a\pi)}.$
This concludes the proof. ∎
Note that (26) has two interesting implications: first, it allows to extend, in an asymptotic perspective, the formula (22) to a large family of segments in the class (C2). Further, it shows that for $a<a^{*}$ the conditioning plateaus around the optimal value, in the sense of the lower bound (23). A numerical verification of Theorem 3.13 is given in Figure 8.
Figure 8: Bounded conditioning of the Vandermonde matrix in the Chebyshev basis for the class (C2), for different values of $a$. Observe that, for $a=0.7$ (orange line in both panels), for $d=6$ there is an exchange in the relationship between $W_{d}]_{1,1}$ and $W_{d}]_{d,d}$.
Remark 3.14.
The situation pictured in Theorem 3.13 is asymptotic. In practice, for intermediate values of $d$, the first case of Proposition 3.11 may happen. Likewise, it may happen that, if $a$ is fixed, there exists a finite value $d^{*}$ for which the largest (respectively the smallest) eigenvalue of $W_{d^{*}}$ exchanges its role with $W_{d^{*}}]_{d^{*},d^{*}}$. This value $d^{*}$ gives the peculiar step pictured in the line corresponding to $a=0.7$ in Figure 8. A detail of this phenomenon is pictured at the right hand panel of the same Figure.
3.2.2 The effect of the Chebyshev basis on generic segments
@@
3.2.3 The normalized case
For the class (C2), the normalization factor $|s_{i}|=|\cos(\tau_{i}-\rho)-\cos(\tau_{i}+\rho)|=2|\sin(\tau_{i})\sin(\rho)|$ turns out to be useful in the computation of the Lebesgue constants performed in [6]. Supposing that $\tau_{i}$ and $\rho$ satisfy the hypotheses of the above section, so that $\sin(\tau_{i})>0$ and $\sin(\rho)>0$, the corresponding Vandermonde matrix is
$\left[\widetilde{V}_{d}\right]_{i,j}=\frac{\frac{2}{j}\sin(j\tau_{i})\sin(j\rho)}{2\sin(\tau_{i})\sin(\rho)}=\frac{\sin(j\tau_{i})\sin(j\rho)}{j\sin(\tau_{i})\sin(\rho)}.$ (27)
Let us observe that $\widetilde{V}_{d}=S_{d}V_{d}$, where $S_{d}$ is the diagonal matrix
$S_{d}]_{i,i}=\frac{1}{2\sin(\tau_{i})\sin(\rho)}$
and $V_{d}$ is the non-normalized matrix (6). Using Binet’s Theorem and Corollary 3.2, we may hence provide a closed formula for $\det\widetilde{V}_{d}$.
Proposition 3.15.
Let $0<\rho<\pi/d$. One has
$\det\widetilde{V}_{d}=\frac{\det V_{d}}{2\sin^{d}(\rho)},$
where a formula for the modulus of $\det V_{d}$ is given in (15), and the sign is $(-1)^{\frac{d(d-1)}{2}}$.
Proof.
It is sufficient to compute $\det S_{d}=\left(\prod_{i=1}^{d}2\sin(\tau_{i})\sin(\rho)\right)^{-1}$. For, we compute
$\prod_{i=1}^{d}2\sin(\tau_{i})\sin(\rho)=\left(\prod_{i=1}^{d}2\right)\left(\prod_{i=1}^{d}\sin(\tau_{i})\right)\left(\prod_{i=1}^{d}\sin(\rho)\right)=2^{d}\sin^{d}(\rho)\left(\prod_{i=1}^{d}\sin(\tau_{i})\right).$
Applying the product formula $\sin(dx)=2^{d-1}\prod_{i=0}^{d-1}\sin\left(\pi i/d+x\right)$ to Chebyshev nodes, and shifting indices, one readily deduces that $\prod_{i=1}^{d}\sin(\tau_{i})=2^{1-d}$, whence
$\det S_{d}=\frac{1}{2^{d}2^{1-d}\sin^{d}(\rho)}=\frac{1}{2\sin^{d}(\rho)}.$
Binet’s Theorem immediately implies the claim. ∎
When $\rho=\pi/(2d)$, one may provide an asymptotics for $\det\widetilde{V}_{d}$. It follows immediately from Corollary 3.2 and the limit $\lim_{d\to\infty}\sin^{d}(\pi/(2d))=(\pi/(2d))^{d}$.
Corollary 3.16.
Let $\rho=\pi/(2d)$. One has
$\lim_{d\to\infty}\left|\det\widetilde{V}_{d}\right|=\frac{\sqrt{d^{3d+1}2^{d-4}}}{d!\pi^{d}}.$
Of course, due to the additional left-multiplication by $S_{d}$, $\widetilde{V}_{d}$ does not satisfy anymore the orthogonality condition of Proposition 3.1. Nevertheless, the conditioning $\kappa_{2}(\widetilde{V}_{d})$ does not grow faster than linearly.
Proposition 3.17.
For every $\rho\in(0,\pi/d)$, one has
$\lim\kappa_{2}(\widetilde{V}_{d})\leq\frac{2\kappa_{2}(V_{d})}{\pi}d.$
Proof.
Let us write $\widetilde{V}_{d}=S_{d}V_{d}$ and analyze $\kappa_{2}(S_{d})$. Clearly, factorizing the term $2\sin(\rho)$, one finds
$\kappa_{2}(S_{d})=\frac{\max_{i}\sin(\tau_{i})}{\min_{i}\sin(\tau_{i})}\leq\frac{1}{\min_{i}\sin(\tau_{i})}.$
The minimum of $\sin(\tau_{i})$ is reached for $\tau_{1}=\pi/(2d)$, the closest to zero Chebyshev node, and for $\tau_{d}=(2d-1)\pi/(2d)$. Considering $\tau_{1}$, as $d\to\infty$, one has $\sin(\pi/2d)\approx\pi/2d$, which implies that
$\kappa_{2}(S_{d})\leq\frac{2d}{\pi}.$
Since $\kappa_{2}(\widetilde{V}_{d})\leq\kappa_{2}(S_{d})\kappa_{2}(V_{d})$, the result immediately follows from Theorem 3.13. ∎
Notice that, for $\rho=\frac{\pi}{2d}$, which corresponds to the parameter $a=1/2$, the upper bound established by Proposition 3.17 simply becomes $\kappa_{2}(\widetilde{V}_{d})\leq d$. The behavior of the conditioning is confirmed by the numerical trend depicted in Figure 10.
Figure 10: Linear conditioning of the Vandermonde matrix in the Chebyshev basis for the class (C2). The parameter is $a=1/2$, so $\kappa_{2}(\widetilde{V}_{d})\leq d$. The resulting segments also belong to the class (C1).